The width of a rectangle is 3 less than half the length. Find the length if the perimeter is 54.

Answer by macston(5194)

Length=L Width=W=1/2L-3
P=2(L+W)Substitute above for W
54 m=2(L+1/2L-3)
54 m=2(1.5L-3)
54 m=3L-6 add 6 to each side
54+6 m=3L-6+6
60=3L divide each side by 3
60/3 m=3L/3
20 m=L ANSWER length=20 meters
CHECK:
W=1/2L-3=1/2(20)-3=10-3=7 m
P=2(L+W)
54 m=2(20m + 7m)
54m=2(27m)
54m=54m